# Kartik Narayan, Vibashan VS, Vishal M. Patel

### v2

@@ 6 Discussion and Future Directions
Table 3: (a) Change in performance of select models under different evaluation settings. (b) Performance improvement seen by leveraging tool-use. (c) Results of finetune experiments on LLaVA1.5 [49] using different data composition.
7 Conclusion
@@
Appendix
@@
In the appendix, we provide a discussion of the limitations of our work. Along with this, we present detailed information regarding the benchmark, including the broad categories, the tasks, the source datasets used, and the dataset statistics. Additionally, we focus on its implementation and provide extensive details about the prompts used for dataset collection and the evaluation strategy. Furthermore, we expand on the results presented in the main paper, providing more details about the baselines, showcasing failure cases of the models, and concluding with an ethics statement.
[…]
1Table of Contents
Appendix A Motivation
@@
In the main paper, we already discussed “Why should MLLMs be proficient in face understanding?”. To reiterate, MLLMs are increasingly deployed as central processors in various advanced applications, including virtual-reality headsets [39], embodied AI [30, 21], driving safety [32, 92], authentication [19], human-computer interaction [103], and sports analysis [109]. In these applications, accurate face understanding is crucial, as face images appear frequently and require accurate face understanding for appropriate responses.
@@
A common question that arises is “Why do we need MLLMs to accomplish face analysis task, which are mostly perception-based, when MLLMs are not explicitly trained for these tasks?”. To clarify, our objective is not to solve face analysis tasks using MLLMs, nor do we aim to outperform specialized face analysis models on established benchmarks. Instead, our focus is on evaluating MLLMs’ ability to understand faces, which is a fundamental aspect of general visual and multimodal reasoning. Strong face understanding is essential for a variety of real-world applications, as discussed in the previous question. Given that MLLMs are expected to serve as general-purpose vision-language models, assessing their performance in this domain provides valuable insights into their limitations and potential areas for improvement. FaceXBench serves as a standardized benchmark to systematically evaluate and monitor the progress of MLLMs in face understanding. By identifying gaps and shortcomings in their capabilities, our work can help guide future advancements in MLLMs, ensuring that they develop stronger and more reliable face understanding skills.
@@
Another important question that comes to mind is “Face analysis tasks are primarily fine-grained and vision-centric, requiring little to no language reasoning. Why is it necessary to evaluate MLLMs on face understanding rather than using specialized vision models for face analysis and then passing the results to LLMs?”. We acknowledge that using an external tool or agent leveraging specialized vision models and then passing the results to LLMs would likely yield better performance in face analysis tasks. To validate this, we have included tool-use results in Table 3(b) of the main paper. However, the primary goal of FaceXBench is not to optimize performance through external assistance but to directly assess the inherent face-understanding capabilities of existing MLLMs. As discussed in the introduction (lines 46–72) of main paper, we argue that MLLMs should possess strong face-understanding skills intrinsically, as this ability is crucial for a wide range of real-world multimodal applications. Many general tasks such as identity-aware reasoning, social interaction analysis, and facial emotion recognition require a model to process and reason about facial features without relying on a separate vision module. Evaluating MLLMs independently allows us to benchmark their progress, identify limitations, and drive improvements in their multimodal reasoning capabilities, ultimately making them more robust for downstream applications where face understanding is essential.
@@
We feel that it also important to answer the question “What is the rationale behind our selected tasks? For example, why do we not include face detection or prompt MLLMs to generate face parsing maps?”. We select tasks that are most relevant to the capabilities and real-world applications of MLLMs. While face detection and direct face segmentation are important computer vision tasks, we exclude them from our benchmark because current MLLMs generally fail to handle them effectively and falls outside the intended use of MLLMs. Using a large MLLM to carry out face detection or parsing is thus impractical , instead, we include tasks such as crowd counting and evaluating segmented attributes because these tests still probe an MLLM’s ability to localize and reason about faces, which is essential for many downstream applications that MLLMs are designed to solve. Again, we would like to emphasize that our goal is not to solve face analysis tasks outright; rather, we select and adapt tasks that suit MLLMs and align with their purpose and intended use.
Appendix B Limitations
@@
Appendix C FaceXBench
In this section of the appendix, we will provide more information on the broad categories and tasks included in the FaceXBench. Additionally, we will provide details on the source datasets and the dataset statistics, with examples of images used in the benchmark.
C.1 FaceXBench Categories
@@ C.3 Source Datasets
C.4 Dataset Statistics
The FaceXBench dataset is derived from 25 public datasets and one newly created dataset. The number of questions sourced from each dataset, along with the type of questions (multiple images, single images, or text-only), as well as the associated tasks and categories, are summarized in Table C.1.
@@ C.4 Dataset Statistics
WMCA [25] 250 100 150 0 Face Anti-spoofing Face Authentication
@@
Table C.1: Question distribution of FaceXBench across datasets
C.5 Images used in the dataset
Figure C.1 displays a subset of the facial images used in the dataset, highlighting the diversity of faces included in the benchmark. The benchmark consists of face images with varying backgrounds, resolutions, and head pose orientations, as well as a wide range of facial expressions. It includes individuals from different age groups, genders, and racial backgrounds, with each face characterized by a versatile set of attributes.
Figure C.1: Collage of a subset of images from the dataset, showcasing the diversity of images used in FaceXBench.
Appendix D Dataset Collection
@@ D.1 Question Templates
D.2 Generating Distractor Options
@@
D.3 FaceXAPI dataset prompt
@@
D.4 Option Extraction from Prediction
  {listing}
@@ D.5 Dataset Format
D.6 Other Evaluation Settings
@@
D.7 Implementation Details
  • VLMEvalKit: We utilized the official open-source implementation, available at https://github.com/open-compass/VLMEvalKit.
Appendix E Results
@@ E.1 Baselines
Idefics2-8b [42] SigLIP-So400m Mistral-7B-v0.1 7B
@@
Table E.1: Vision encoder, LLM and LLM size of the baseline models
E.2 Results Across Tasks
To provide additional insights, we summarize the task-wise performance of various models in Table E.2. We observe that models struggle with tasks such as head pose estimation, crowd counting, and deepfake detection. Conversely, tasks like gender prediction, facial expression recognition, and high-resolution face recognition are comparatively easier. Additionally, we present the results of selected models under the ”in-context” and ”chain-of-thought” evaluation settings in Table E.4 and Table E.4, respectively. We observe a consistent drop in performance, indicating that the models are not capable of effectively processing in-context information and fail to reason in tasks related to face understanding.
Models Overall
Expression
Recognition
Age
Estimation
Race
Estimation
Crowd
Counting
Celebrity
Identification
HR FR LR FR
Face
Anti-spoofing
Tools
Retrieval
Attributes
Prediction
Gender
Prediction
Headpose
Estimation
Face
Parsing
Deepfake
Detection
Open source MLLMs ($\mathbf{<4}\text{B}$ parameters)
Open source MLLMs ($\mathbf{4}\text{B}$ $\mathbf{13}\text{B}$ parameters)
Open source MLLMs ($\mathbf{>13}\text{B}$ parameters)
Proprietary MLLMs
Table E.2: Performance of various models across all evaluation tasks of FaceXBench.
@@ E.2 Results Across Tasks
Table E.3: (a) Performance of models in in-context evaluation setting.
@@
Table E.4: (b) Performance of models in chain-of-thought evaluation setting.
E.3 Human and SOTA Vision models evaluation details
@@
As can be seen from Table E.2, there are several tasks where human participants perform significantly worse than SOTA models, primarily due to the nature of the tasks. These include deepfake detection, head pose estimation, celebrity identification, crowd counting, and age estimation. In deepfake detection, some manipulations are more pronounced in the spectral domain but remain imperceptible to the human eye. Similarly, age estimation requires high precision, as the distractor options are designed to be very close in value, meaning even a small human error results in an incorrect choice, making the benchmark more challenging and forcing the model to be precise. For head pose estimation, humans struggle to visualize fine angular changes in yaw, pitch, and roll, leading to poor performance. Additionally, the low human performance in celebrity identification is primarily due to the participants’ limited familiarity with a wide range of celebrities.
E.4 Failure Cases
We conducted a failure case analysis using top-performing models, namely GPT-4o, Qwen2-VL, and InternVL2. As part of this analysis, we present one question from each task to illustrate the failure cases. Our observations reveal that these models struggle to answer basic questions related to face understanding, highlighting the need for a dedicated benchmark to monitor the progress of MLLMs in this domain.
@@ E.4 Failure Cases
Face Anti-spoofing Question: <image 1><image 2> Which statement best describes the images? Option: (A) Image 1 is bonafide and Image 2 is attack (B) Image 1 is attack and Image 2 is bonafide (C) Both images are bonafide (D) Both images are attacks [Uncaptioned image] [Uncaptioned image] <image 1> <image 2> GPT-4o: B Prediction: Qwen2-VL: B Prediction: InternVL2: (B) Image 1 is attack and Image 2 is bonafide Prediction: Ground Truth: (D) Both images are attacks
@@
Appendix F Ethical Considerations

### v3

@@ V Results
TABLE IV: (a) Change in performance of select models under different evaluation settings. (b) Performance improvement seen by leveraging tool-use. (c) Results of finetune experiments on LLaVA1.5 [46] using different data composition.
Results across different evaluation categories. The performance of various models on FaceXBench, as shown in Table II highlights the limited face understanding capabilities of current models. Specifically, they perform poorly on tasks such as face authentication and face localization, which require fine-grained facial feature extraction and spatial understanding of facial structure. The low performance in the “Bias & Fairness” category suggests that existing MLLMs exhibit biases toward certain age, gender, and racial groups, which need to be mitigated prior to deployment. Surprisingly, GPT-4o performs poorly in the “Bias & Fairness category”, but on examination, it was found that the model often chose not to answer due to its safety alignment. The top-performing models achieve around 7070% performance in face recognition tasks; however, their performance drops significantly on low-resolution face recognition. This decline can be attributed to training data predominantly comprising of high-resolution images, which limits models’ effectiveness on low-resolution inputs. Some models may have been trained on image-text datasets with face images, such as FFHQ-Text [137], CelebA-Dialog [34], LAION-Face [136], FaceCaption-15M [18], as part of large-scale pretraining. Although these datasets often contain attribute and expression information in text, the models still perform poorly in the “Face Analysis” section, indicating substantial room for improvement.
Open Source vs Closed Source Models. We observe that open-source models, such as InternVL2 and Qwen2-VL, outperform proprietary models like GPT-4o and GeminiPro 1.5, achieving accuracies of 57.8057.80% and 57.8657.86%, compared to 50.5050.50% and 56.9656.96%, respectively. Generally, MLLMs have shown a trend where closed-source models outperform open-source ones. However, in the sensitive domain of face analysis, proprietary models are safety-aligned before deployment. The relatively poor performance of proprietary models in “Bias & Fairness” and “Face Analysis” is primarily due to content moderation constraints. Notably, GeminiPro 1.5 demonstrates comparatively better performance than other models on “Face Tools Use”, showcasing its ability to leverage specialized tools for complex scenarios involving multiple face-related tasks.
Performance across various tasks. To gauge the difficulty of various tasks, we plot the performance of the top-5 models (44B-1313B parameters) across each task, as shown in Figure 4(a). To provide task-specific insights, we summarize the task-wise performance of various models in Table III. We observe that, on average, models struggle with tasks such as crowd counting, deepfake detection, head pose estimation, and low-resolution face recognition, highlighting areas where MLLMs’ need improvement. In contrast, gender prediction is one of the easier tasks, with an average performance of approximately 8080%. Additionally, we plot the average performance of these top-5 models on multiple-image and single-image question. Figure 4(c) shows that models generally perform poorly on questions with multiple images as inputs compared to single-image questions. This is primarily because the models are required to process more visual information and compare and contrast facial features across multiple images, making the questions more challenging. Furthermore, models such as Qwen2-VL [106] and LLaVA-OneVision [46], which use dynamic resolution to handle arbitrary image resolutions, perform better than models on segmentation task. This mapping of images to a dynamic number of visual tokens more closely resembles human-like visual processing and leads to improved performance in face-understanding tasks.
Effect of LLM and it’s size on performance. To analyze the impact of LLM on model performance, we plot a performance curve for models that share the same SigLIP SO400M/14@384 vision encoder but differ in their LLM backbone. From Figure 4(b), we observe that performance improves with an increase in LLM size. Furthermore, when comparing LLMs of different sizes within the same family, such as Qwen2, we see that the curve shape remains consistent, with performance values translating upward across all dimensions. This indicates that as LLM size increases, model capabilities improve proportionately across various dimensions, maintaining a similar pattern of performance enhancement.
Performance under different evaluation settings. We evaluate three selected models under the in-context task description and chain-of-thought settings, with results summarized in Table. IV. In the in-context setting, we observe some performance improvements in face authentication and face tools use. However, overall performance drops, indicating that the models struggle to utilize in-context information effectively. In the chain-of-thought setting, we observe a substantial decline in performance, suggesting that, although these models exhibit reasoning capabilities, they do not transfer effectively to face understanding.
VI Discussion and Future Directions
In this section, we discuss possible future directions to improve the face understanding capabilities of MLLMs. The MLLM community has developed numerous supervised fine-tuning (SFT) datasets, such as COCO Caption [11], ScienceQA [84], Vision-FLAN [115], ChartQA [66], FigureQA [36], Geometry3K [24], MAVIS MCollect [127], MATHQA [6], TextOCR [91], OCR-VQA [67], and MagpiePro [113]. These supervised fine-tuning sets target various skills, including document and chart understanding, mathematics, fine-grained perception, grounding, reasoning, and general OCR. Developing these general skills helps models perform better on existing tasks and enhances their reasoning and visual processing capabilities for improved zero-shot task performance.
However, to answer the question, “Will supervised fine-tuning improve the performance of MLLMs on face understanding?” we conducted multiple experiments by finetuning LLaVA1.5 [57] on a random 7070k subset of the FaceCaption15M [18], which comprises image-text pairs containing attribute, age, and gender information. We fine-tuned the vision projector and the LLM backbone using LoRA [28] and summarized the results for various data compositions in Table IV(c). Our findings reveal that naive fine-tuning on MLLM with only Face SFT data results in poor performance (Table IV(c), row 1), as models tend to lose their general reasoning and perception capabilities. We incorporated the complete LLaVA SFT dataset (665665k samples) alongside the Face SFT, and the model achieved a score of 35.18 (Table IV(c), row 2). We randomly sampled 200200k examples from the full 665665k LLaVA 1.5 dataset, combining them with the 70k Face SFT data for fine-tuning, resulting in an improved score of $35.24$ (Table IV(c), row 3). This experiment demonstrates the benefit of integrating FaceSFT data in the right proportion with existing reasoning and instruction-tuning datasets.
We explored an alternative research direction to improve MLLMs’ performance on face understanding using specialized tools. To investigate the benefits of tool use, we selected specific datasets and converted predictions from state-of-the-art models into text, which we then provided as context to the MLLM. As shown in Table IV(b), this approach resulted in a significant performance boost. We believe equipping MLLMs with agentic behavior through specialized tools is a promising way to enhance their face understanding capabilities. For Future Work: We advocate the development of diverse supervised fine-tuning sets covering various aspects of face understanding and training MLLMs to leverage specialized tools. Following this direction, we believe future MLLMs will be capable of advanced face understanding, with FaceXBench serving as a critical resource for monitoring progress by benchmarking models across multiple dimensions of face understanding.
VII Conclusion
@@
Acknowledgment
This research is based upon work supported in part by the Office of the Director of National Intelligence (ODNI), Intelligence Advanced Research Projects Activity (IARPA), via [2022-21102100005]. The views and conclusions contained herein are those of the authors and should not be interpreted as necessarily representing the official policies, either expressed or implied, of ODNI, IARPA, or the U.S. Government. The US Government is authorized to reproduce and distribute reprints for governmental purposes notwithstanding any copyright annotation therein.
Supplementary Material
@@ Acknowledgment
In the supplementary material, we discuss additional questions that frame the motivation for our work and provide a discussion of its limitations. Along with this, we present detailed information regarding the benchmark, including the broad categories, the tasks, the source datasets used, and the dataset statistics. Additionally, we focus on its implementation and provide extensive details about the prompts used for dataset collection and the evaluation strategy. Furthermore, we expand on the results presented in the main paper, providing more details about the baselines, showcasing failure cases of the models, and concluding with an ethics statement.
VIII Motivation
@@
We already discussed “Why should MLLMs be proficient in face understanding?”in Introduction I. To reiterate, MLLMs are increasingly deployed as central processors in various advanced applications, including virtual-reality headsets [39], embodied AI [30, 21], driving safety [32, 92], authentication [19], human-computer interaction [103], and sports analysis [109]. In these applications, accurate face understanding is crucial, as face images appear frequently and require accurate face understanding for appropriate responses. Below we answer some additional questions that frame our motivation and benchmark design.
@@
Why do we need MLLMs to accomplish face analysis task, which are mostly perception-based, when MLLMs are not explicitly trained for these tasks?”. To clarify, our objective is not to solve face analysis tasks using MLLMs, nor do we aim to outperform specialized face analysis models on established benchmarks. Instead, our focus is on evaluating MLLMs’ ability to understand faces, which is a fundamental aspect of general visual and multimodal reasoning. Strong face understanding is essential for a variety of real-world applications, as discussed in the previous question. Given that MLLMs are expected to serve as general-purpose vision-language models, assessing their performance in this domain provides valuable insights into their limitations and potential areas for improvement. FaceXBench serves as a standardized benchmark to systematically evaluate and monitor the progress of MLLMs in face understanding. By identifying gaps and shortcomings in their capabilities, our work can help guide future advancements in MLLMs, ensuring that they develop stronger and more reliable face understanding skills.
@@
Face analysis tasks are primarily fine-grained and vision-centric, requiring little to no language reasoning. Why is it necessary to evaluate MLLMs on face understanding rather than using specialized vision models for face analysis and then passing the results to LLMs?”. We acknowledge that using an external tool or agent leveraging specialized vision models and then passing the results to LLMs would likely yield better performance in face analysis tasks. To validate this, we have included tool-use results in Table 3(b) of the main paper. However, the primary goal of FaceXBench is not to optimize performance through external assistance but to assess the inherent face-understanding capabilities of existing MLLMs. As discussed in the introduction (lines 46–72) of main paper, we argue that MLLMs should possess strong face-understanding skills intrinsically, as this ability is crucial for a wide range of real-world multimodal applications. Many general tasks such as identity-aware reasoning, social interaction analysis, and facial emotion recognition require a model to process and reason about facial features without relying on a separate vision module. Evaluating MLLMs independently allows us to benchmark their progress, identify limitations, and drive improvements in their multimodal reasoning capabilities, ultimately making them more robust for downstream applications where face understanding is essential.
@@
What is the rationale behind our selected tasks? For example, why do we not include face detection or prompt MLLMs to generate face parsing maps?”. We select tasks that are most relevant to the capabilities and real-world applications of MLLMs. While face detection and direct face segmentation are important computer vision tasks, we exclude them from our benchmark because current MLLMs generally fail to handle them effectively and falls outside the intended use of MLLMs. Using a large MLLM to carry out face detection or parsing is thus impractical , instead, we include tasks such as crowd counting and evaluating segmented attributes because these tests still probe an MLLM’s ability to localize and reason about faces, which is essential for many downstream applications that MLLMs are designed to solve. Again, we would like to emphasize that our goal is not to solve face analysis tasks outright; rather, we select and adapt tasks that suit MLLMs and align with their purpose and intended use.
Why we not include open-ended visual question answering in the benchmark ?”. We chose not to include open-ended visual question answering (VQA) in our benchmark, following the precedent set by leading benchmarks such as MMBench [61] [ECCV 2024], SEEDBench [47] [CVPR 2024], MMMU [122] [CVPR 2024], and MathVista [64] [ICLR 2024], which rely exclusively on multiple-choice questions (MCQs). The MCQ format offers key advantages: it enables standardized, reproducible, and scalable evaluation, while eliminating the ambiguity and subjectivity inherent in open-ended answers, especially critical in the facial domain, where expressions and attributes can be nuanced and difficult to judge consistently. In contrast, open-ended VQA often depends on human evaluation or external matching models, introducing non-determinism and inconsistency. By using MCQs, we ensure deterministic ground truth and fair comparisons across models. Importantly, MCQs do not simplify the reasoning task; models must still perform complex interpretation and decision-making. The format merely standardizes the output, making the evaluation process more reliable and accessible, and helping drive community-wide adoption and progress in facial understanding tasks.
IX Limitations
@@
X FaceXBench
In this section of the supplementary material, we will provide more information on the broad categories and tasks included in the FaceXBench. Additionally, we will provide details on the source datasets and the dataset statistics, with examples of images used in the benchmark.
X-A FaceXBench Categories
@@ X-C Source Datasets
X-D Dataset Statistics
The FaceXBench dataset is derived from 25 public datasets and one newly created dataset. The number of questions sourced from each dataset, along with the type of questions (multiple images, single images, or text-only), as well as the associated tasks and categories, are summarized in Table X.1.
@@ X-D Dataset Statistics
WMCA [25] 250 250 0 0 Face Anti-spoofing Face Authentication
@@
TABLE X.1: Question distribution of FaceXBench across datasets
X-E Images used in the dataset
Figure X.1 displays a subset of the facial images used in the dataset, highlighting the diversity of faces included in the benchmark. The benchmark consists of face images with varying backgrounds, resolutions, and head pose orientations, as well as a wide range of facial expressions. It includes individuals from different age groups, genders, and racial backgrounds, with each face characterized by a versatile set of attributes.
Figure X.1: Collage of a subset of images from the dataset, showcasing the diversity of images used in FaceXBench.
XI Dataset Collection
@@ XI-A Question Templates
XI-B Generating Distractor Options
@@
XI-C FaceXAPI dataset prompt
@@
XI-D Option Extraction from Prediction
 {listing*}
@@ XI-E Dataset Format
XI-F Other Evaluation Settings
@@
XII Results
@@ XII-A Baselines
Idefics2-8b [laurençon2024matters] SigLIP-So400m Mistral-7B-v0.1 7B
@@
TABLE XII.1: Vision encoder, LLM and LLM size of the baseline models
XII-B Results across different evaluation settings
To provide additional insights, we present the results of selected models under the "in-context" and "chain-of-thought" evaluation settings in Table XII.3 and Table XII.3, respectively. We observe a consistent drop in performance, indicating that the models are not capable of effectively processing in-context information and fail to reason in tasks related to face understanding.
@@ XII-B Results across different evaluation settings
TABLE XII.2: (a) Performance of models in in-context evaluation setting.
@@
TABLE XII.3: (b) Performance of models in chain-of-thought evaluation setting.
XII-C Human and SOTA Vision models evaluation details
@@
As can be seen from Table III, there are several tasks where human participants perform significantly worse than SOTA models, primarily due to the nature of the tasks. These include deepfake detection, head pose estimation, celebrity identification, crowd counting, and age estimation. In deepfake detection, some manipulations are more pronounced in the spectral domain but remain imperceptible to the human eye. Similarly, age estimation requires high precision, as the distractor options are designed to be very close in value, meaning even a small human error results in an incorrect choice, making the benchmark more challenging and forcing the model to be precise. For head pose estimation, humans struggle to visualize fine angular changes in yaw, pitch, and roll, leading to poor performance. Additionally, the low human performance in celebrity identification is primarily due to the participants’ limited familiarity with a wide range of celebrities.
XII-D Failure Cases
We showed one example in the main paper. We have included examples for the other tasks below. Our observations reveal that these models struggle to answer basic questions related to face understanding, highlighting the need for a dedicated benchmark to monitor the progress of MLLMs in this domain.
@@ XII-D Failure Cases
XIII Ethical Considerations